OpenStudy (anonymous):

tan ( cos^-1 4/5 + tan^-1 2/3) = ?

OpenStudy (shubhamsrg):

 any help ?

OpenStudy (shubhamsrg):

i mean does this help * ?

OpenStudy (anonymous):

cos^-1 4/5 = 37

OpenStudy (shubhamsrg):

you didnt need to know that for this ques..

hartnn (hartnn):

maybe......use first tan (A+B),then tan(cos^-1 x) = sqrt(1-x^2)/x

OpenStudy (aravindg):

cos^-1 4/5=tan^-13/4

OpenStudy (aravindg):

hence the qn becomes tan(tan^-13/4+tan^-1 2/3)

OpenStudy (shubhamsrg):

yes you may do that,, what i was trying to state is from the fig, or otherwise , cos^-1 4/5 = tan^-1 3/4 on that subsn, use identity of tan^-1 a + tan^-1 b then its quite simple..

OpenStudy (shubhamsrg):

as @AravindG has stated already..

OpenStudy (anonymous):

hw cos^-1 4/5=tan^-13/4 ???